What is the molarity of a solution containing 32g of K2SO4 IN 350 cm3 of solution?
1
Expert's answer
2020-12-06T13:57:06-0500

molarity: c=n/V

n(K2SO4)=m/M=32g / 174g/mol = 0.184mol

V=350cm3=0.35L

c(K2SO4)=0.184mol / 0.35L = 0.526 mol/L